LifewithJayda10:

. Find the percent of discount. Round to the nearest tenth. original price $24,365; selling price $16,820

destineyminter2005:

have you learned how to find the discount price

LifewithJayda10:

no

destineyminter2005:

24365x=16820

destineyminter2005:

x=24365/16820

destineyminter2005:

so what answer do you get

destineyminter2005:

@destineyminter2005 wrote:
x=24365/16820
reverse the oder of those sorry

LifewithJayda10:

can u divide it for me

destineyminter2005:

0.69 and now round it to the nearest tenths place

destineyminter2005:

so since the number to the right of 6 is greater than 5 you round up so it would be 7.0

destineyminter2005:

@LifewithJayda10 did that help

Florisalreadytaken:

you can use the rule of three to simply find the percentage difference, or the discount: \[ \frac{$24365}{100\%} = \frac{$16820}{x \%} \] \[ $24365x = $16820 \times 100 \%\] \[ $24365x = 1682000 \frac{$}{\%} \] \[ x=\frac{ 1682000 \frac{\cancel{$}}{\%} }{\cancel{$}24365} \] \[ x \approx 69\% \] so \( 69 \% \) would be how much the 2nd price is -- thus, to find the discount you do this: \[ 100\%-69\%=\color{lightskyblue}{31\%} \] so that would be your answer

jhonyy9:

@destineyminter2005 wrote:
so since the number to the right of 6 is greater than 5 you round up so it would be 7.0
i think will be 0,7 not 7,0 ATTENTION !!!

destineyminter2005:

@jhonyy9 wrote:
@destineyminter2005 wrote:
so since the number to the right of 6 is greater than 5 you round up so it would be 7.0
i think will be 0,7 not 7,0 ATTENTION !!!
yes im sorry i posted it wrong ok